OpenStudy (anonymous):

lim (h->0) cos(pi/2 +h) - cos (pi/2) /h Using definition of a derivative

OpenStudy (anonymous):

\[\cos ((\pi/2) + h) - \cos (\pi/2) /h \] this looks a little cleaner